OpenStudy (anonymous):

is this additive inverse property ? 1 + 2 + 3 = 1 + (2 + 3) is this additive inverse property ? 1 + 2 + 3 = 1 + (2 + 3) @Mathematics

OpenStudy (asnaseer):

a = a is the identity statement a+0 = a tells us that zero is the additive identity a*1 = a tells us that one is the multiplicative identity a+b = b+a is the commutative property of addition ab = ba is the commutative property of multiplication a+(b+c) = (a+b)+c is the associative property of addition a(bc) = (ab)c is the associative property of multiplication